OpenStudy (anonymous):

slope of line that passes through (-2,8) with slope of 2

hero (hero):

So they already give you the slope m = 2, but you still have to find the slope anyway?

OpenStudy (anonymous):

i worded tht wrong........standard form equation of the line passing through (-2, 8) with a slope of 2?

hero (hero):

Plug in given values into y = mx + b, then find b: 8 = 2(-2) + b 8 = -4 + b 8 - 4 = b 4 = b Write the equation of the line with m = 2, b = 4 y = 2x + 4 subtract y from both sides, then subtract 4 from both sides: 2x - y = -4

OpenStudy (anonymous):

thank you